Hi all, please help me to solve this - I have the answer, but I cannot "back engineer" how we arrived at it. Thank you!
The gradient of the line joining (4, q) to (6, 5) is twice the gradient of the line joining (0, 0) to (4, q). Find q.

Answer by stanbon(75887)

The gradient of the line joining (4, q) to (6, 5) is twice the gradient of the line joining (0, 0) to (4, q). Find q.
-----
Equation:
(q-5)/(4-6) = 2[(q-0)/(4-0)]
-----
(q-5)/(-2) = 2[q/4]
q-5 = -4(q/4)
4q-20 = -4q
8q = 20
q = 2.5
-------------
Cheers,
Stan H.
------------